5. Find the absolute max and absolute min of f(:1, y) = x2 + 2y2 – 2.0 – 4y on the rectangle (<r<2,0 <y<3.
Answer #1

(3) Cu -y=3 x=0 x=2 5) Giren foc, y) = x²+242_2x-44 99 (213) € Osxsa; osys3 critical points occurs when [fc zot yo coros (210points & on line yzo foo = x² to-2x -0 =x²-2x f(0) = 2x-2 f(x)=0 2x-220 (2=D point is Clo) on line y=3 F(x) = x² +2(3)² - 2x